# Technical report on label-informed logit redistribution for better domain generalization in low-shot classification with foundation models

### v2

@@ 4.1 Experimental Setup
Datasets. We evaluated it our method RccPL on $12$ standard vision benchmark datasets that includes, Flowers102 Nilsback and Zisserman (2008), StanfordCars Krause et al. (2013), Imagenet Deng et al. (2009), DTD Cimpoi et al. (2014), SUN397 Xiao et al. (2010), CIFAR10101010 Coates et al. (2011), EuroSAT Helber et al. (2019), OxfordPets Parkhi et al. (2012), UCF101 Soomro (2012), Food101 Bossard et al. (2014), FGVCAircraft Maji et al. (2013), Caltech101 Fei-Fei et al. (2004). For domain generalisation we used ImageNet-1K with its four variants like ImageNet A (Art) Hendrycks et al. (2021a), ImageNet-V2 (V2) Recht et al. (2019), ImageNet R (Real) Hendrycks et al. (2021b) and ImageNet S (Sketch) Wang et al. (2019) to ensure its performance.
Implementation details. Our experiment is in two parts. In first part we evaluate RccPL performance for prompt learning methods. To achieve this, We used pre-trained CLIP (ViT-B16) Radford et al. (2021) model in prompt learning experiments. In second part, we use CLIP (ViT-B16) and (ResNet-50) as backbone while comparing RccPL performance with CLIPCalib Murugesan et al. (2025)  [^1] The CLIPCalib prompt learning comparison results are reproduced by using their code published in paper which is available at https://github.com/Bala93/CLIPCalib. in domain generalization. To check the performance of RccPL we performed following sets of experiments.
@@
  • CMP integration enhances calibration in domain generalization. To analyze the robustness of RccPL, we performed experiment in domain generalization settings. The results are presented in tabular formate along with detail discussion in section 5 and appendix B.
  • RccPL improvement against benchmarks. To check the calibration performance of RccPL in comparison to benchmarking, we perform detail experiments with two backbone (ResNet-50 and ViT-B16). The results are discussed in section 5 and appendix B.
Due to spatial constraint more experimental detail can be found in appendix B.
Baselines. We compare RccPL performance with existing prompt learning methods, such as CoOp Zhou et al. (2022a), CoCoOp Zhou et al. (2022b), ProDA Lu et al. (2022), MaPLe Khattak et al. (2023), and KgCoOp Yao et al. (2023).
Benchmarks. We compared RccPL confidence normalization performance against standard CLIP in zero-shot setting, with zero-shot calibration method CLIPCalib Murugesan et al. (2025).
Evaluation metrics. To measure the confidence calibration of RccPL, we used three evaluation metrics Expected Calibration Error (ECE) Guo et al. (2017), Adaptive Calibration Error (ACE) Guo et al. (2017) and Maximum Calibration Error (MCE) Nixon et al. (2019) and accuracy in our work.
Table 1: RccPL average calibration performance across 11 datasets for downstream tasks. the row with + shows our method combined with existing prompt learning method. The $\downarrow$ represent the smaller value is better, while the $\uparrow$ in accuracy (Acc) denotes the larger value is significant. Bold values shows more significant result.
@@
KgCoOp + (RccPL) 76.3 4.11 0.97 4.07
Table 2: Comparison of RccPL when combined with prompt learning methods. Bold values represent significant performance. In the left column Acc $\uparrow$ demonstrates larger value is significant. In the right column ECE $\downarrow$ shows smaller value is significant.
@@
KgCoOp + (RccPL) 72.9 74.3 48.8 52.5 65.4 60.3 2.92 4.13 7.08 13.4 3.08 6.92
Table 3: RccPL results in comparison with CLIPCalib Murugesan et al. (2025) with ResNet-50 backbone . RccPL results for ViT-16 backbone implementation are given in the Appendix B. Bold value represents significance improvement in RccPL performance. In ACC part ($\uparrow$ ) denotes larger value better accuracy. while in ECE part ($\downarrow$ ) demonstrate smaller value better calibration.
@@
CLIPRccPL 3.41 3.30 3.52 3.57 3.29 3.36 3.40 3.19 3.46 3.61 3.17 3.31
Table 4: RccPL performance results in comparison with CLIPCalib with (ViT-B16) as backbone for domain genralisation .
@@ 5 Discussions
In our work, we evaluated the effectiveness of CMP calibration using ECE, ACE, MCE and generalization of CMP using acc metrics. But due to space constraint we reported ECE results in this section. The remaining metrics results can be found in B.
Calibration performance of CLIP-based adaptive models. Figure 1 shows confidence calibration of CLIP in zero-shot settings for CIFAR10101010 classification task. The top row (left) in figure shows contrastive loss result without integration of CMP into loss, it has been observed that the confidence distribution for predicting class 00 is spread across all incorrect classes. This indicates overconfidence of the model as the model logits are less focused. The top row (right) shows results of CMP loss for zero-shot tasks, it is shown in figure that the confidence distribution is concentrated near lower values of confidence. This indicates that CMP reduces the overconfidence for incorrect prediction.
@@
The results in Table 1 indicate that integrating CMP consistently reduces ECE, thereby improving model calibration. For example, the baseline CoOp achieves an ECE of $3.16$, while its penalty-enhanced version (CoOp + RccPL) reduces ECE to $2.98$, representing a $5.70\%$ improvement. Similarly, the ProDA baseline shows an ECE of $4.13$, which decreases to $3.94$ with the penalty-based version (ProDA + RccPL), reflecting a $4.60\%$ reduction.
Notably, MaPLe demonstrates the highest improvement, with ECE decreasing from $3.19$ to $2.88$, corresponding to a $9.70\%$ reduction, highlighting the strong impact of CMP for this specific prompt learning method. In the case of KgCoOp, integrating CMP (KgCoOp + RccPL) reduces ECE from $4.24$ to $4.07$, marking a $4.00\%$ improvement.
@@
CMP integration enhances calibration in domain generalization. To evaluate the calibration performance of RccPL integration, we analyze the ECE across domain generalization datasets. The results are reported in Table 2. We discuss the improvement in ECE for each prompt learning method while focusing on source and target domain. It is observed in Table 2, CoOp achieve an average ECE of $8.01$ reflecting significant calibration error. By incorporating our penalty to this method the average ECE is reduced to $7.32$ representing $8.56$% improvement. The ECE improvement shows the effective mitigation of overconfidence in CoOp method, particularly in domain generalizing setting. The reported ECE for CoCoOp without RccPL is $6.83$, while penalized CoCoOp ECE is reduced to $6.15$, yielding $9.8$% improvement in calibration error under challenging target domain.
@@
ProDA baseline achieves $8.10$% ECE on average, with penalty integration ECE reduced to $7.32$ which is $9.60$% improvement. The baseline method MaPLe shows an average of $6.48$ ECE, which is reduced to $5.25$ by integration of our penalty, marking on average $19.0$% performance improvement. KgCoOp shows an average ECE of $7.38$ as baseline, while penalty based KgCoOp reports $6.92$ corresponding $6.20$% reduction in ECE.
@@
From the results given in Table 2, RccPL consistently improves in ECE across domain generalization datasets with improvement within range ($6.20$% - $19.0$%). The results confirms the effectiveness of RccPL in mitigating overconfidence under challenging scenarios of domain generalization settings.
@@
RccPL improvement against benchmarks. RccPL calibration performance evaluation results in comparison with benchmarking method CLIPCalib and CLIP (baseline) for vision datasets and domain generalization datasets are given in Table 3 and Table 4 respectively.
@@
In Table 3, it can be observed that CLIPRccPL achieves an average ECE of $3.31$ significantly improving upon the baseline average ECE i.e $4.49$% and $2.20$% improvement over benchmark CLIPCalib. The calibration improvement is notable on various datasets like Flowers102 ($3.76$ to $3.29$), EuroSAT ($3.42$ to $3.17$), and Aircraft ($3.44$ to $3.19$), where RccPL excels in resolving overconfidence. The results are clearly in favor of our method, indicating the importance of RccPL integration into CLIP framework leads to better-calibrated models. The consistent reduction in ECE across datasets demonstrates the robustness of our method.
@@
Table 4 denotes results of RccPL calibration results in comparison with CLIP (baseline) and CLIPCalib benchmark on domain generalization datasets.
@@
The average calibration error reduces from $5.01$ to $4.71$ when penalty is plugged-in into CLIP framework, showing an improvement of $5.94$% as compared to baseline while $2.79$% over the baseline in challenging datasets. We have noticed that our method consistently outperform the benchmark across domain generalization datasets. We have also reported the result for RccPL by changing the backbone of CLIP which are given in appendix B.
@@
Tuning strength of CMP.The regularization strength $\lambda$ is chosen via model selection using line search. Models are evaluated against ECE as the metric, as shown in appendix, Table 7. We report ECE results within range ($0.001$ $0.950$). We have noticed that ECE score is better for a value of $lambda$ ($\lambda$ = $0.01$). All of our experiment results are reported with same value of $\lambda$. It is notable that small regularization such as this amount is sufficient to enable model confidences to realign themselves with the annotated classification objective. More detail results can be found in Table 7.
@@ 6 Conclusions
Appendix A Theoretical Guarantees
Proposition A.1.
@@ Appendix A Theoretical Guarantees
Let $P(x,y^{\prime})=\frac{\exp(s(x,y^{\prime}))}{\sum_{k}\exp(s(x,y_{k}))}$ is softmax probability for correct pair. By def. of CMP:
$\text{CMP}=\frac{P(x,y^{\prime})}{\sum_{y_{j}\neq y^{\prime},\,P(x,y_{j})>P(x,y^{\prime})}P(x,y_{j})},$
@@
$P(x,y_{j})\leq 1\quad\forall j,\quad\text{and}\quad\sum_{j}P(x,y_{j})=1,$
@@
$P(x,y^{\prime})=\sum_{P(x,y_{j})>P(x,y^{\prime})}P(x,y_{j})\implies\text{CMP}=1$
@@
Proposition A.2.
The final combined CLIP loss $\mathcal{{L}_{\text{final}}}$ has a well-defined lower bound and ensures convergence during optimization.
Proof.
The final combined CLIP loss is given by:
$\mathcal{{L}_{\text{final}}}=\mathcal{L}_{\text{c}}+\lambda\cdot\mathcal{L}_{\text{CMP}},$
For the CLIP contrastive loss $\mathcal{L}_{\text{c}}$, the softmax probabilities ensure:
$-\log P(x,y^{\prime})\geq 0.$
For the CMP penalty term:
$\frac{P(x,y^{\prime})}{\sum_{y_{j}\neq y^{\prime},\,P(x,y_{j})>P(x,y^{\prime})}P(x,y_{j})+\epsilon}\leq 1,$
as $\text{CMP}\leq 1$ by definition, and $\lambda>0$.
Thus, the CMP term does not decrease the overall loss below the contrastive loss $\mathcal{L}_{\text{c}}$. Combining these, the per-sample loss is bounded below by the contrastive loss:
$\mathcal{{L}_{\text{final}}}\geq\mathcal{L}_{\text{c}}.$
Since $\mathcal{L}_{\text{c}}\geq 0$, the combined loss $\mathcal{{L}_{\text{final}}}$ is also bounded below by 00, ensuring stable optimization.
This completes the proof. ∎
Proposition A.3.
@@
$\mathcal{L}_{\text{CMP}}=\frac{1}{N}\sum_{i=1}^{N}\frac{P(x_{i},y_{i})}{\sum_{y_{j}\neq y_{i},\,P(x_{i},y_{j})>P(x_{i},y_{i})}P(x_{i},y_{j})+\epsilon}.$
@@
$P(x,y^{\prime})>P(x,y_{j})\quad\forall y_{j}\neq y^{\prime},$
@@
$\mathcal{L}_{\text{CMP}}(x,y^{\prime})\approx 0\quad\text{for correctly classified pairs.}$.
@@
$P(x,y_{j})>P(x,y^{\prime})\quad\exists(y_{j}\neq y^{\prime}),$
@@
$\mathcal{L}_{\text{CMP}}=\frac{1}{N}\sum_{i=1}^{N}\frac{P(x_{i},y_{i})}{\sum_{y_{j}\neq y_{i},\,P(x_{i},y_{j})>P(x_{i},y_{i})}P(x_{i},y_{j})+\epsilon}.$
@@
$\frac{\partial\mathcal{L}_{\text{CMP}}}{\partial s(x_{i},y_{i})}=\frac{\lambda}{N}\cdot\frac{\partial P(x_{i},y_{i})}{\partial s(x_{i},y_{i})}\cdot\frac{\sum_{P(x_{i},y_{j})>P(x_{i},y_{i})}P(x_{i},y_{j})+\delta-P(x_{i},y_{i})}{\left(\sum_{P(x_{i},y_{j})>P(x_{i},y_{i})}P(x_{i},y_{j})+\delta\right)^{2}}.$
Since $P(x_{i},y_{i})\leq 1$, the gradient is bounded. The term $\delta>0$ here refers to the infinitesimal step size. Furthermore, the gradient decays smoothly as the gap increases between $P(x_{i},y_{i})$ and $\sum_{P(x_{i},y_{j})>P(x_{i},y_{i})}P(x_{i},y_{j})$, ensuring numerical stability.
@@
A.1 Relation to Perplexity

### v3

@@ 4.1 Experimental Setup
Datasets. We evaluated our method CMP on $12$ standard vision benchmark datasets that includes, Flowers102 Nilsback and Zisserman (2008), StanfordCars Krause et al. (2013), Imagenet Deng et al. (2009), DTD Cimpoi et al. (2014), SUN397 Xiao et al. (2010), CIFAR1010 Coates et al. (2011), EuroSAT Helber et al. (2019), OxfordPets Parkhi et al. (2012), UCF101 Soomro (2012), Food101 Bossard et al. (2014), FGVCAircraft Maji et al. (2013), Caltech101 Fei-Fei et al. (2004). For domain generalisation we used ImageNet-1K with its four variants like ImageNet A (Art) Hendrycks et al. (2021a), ImageNet-V2 (V2) Recht et al. (2019), ImageNet R (Real) Hendrycks et al. (2021b) and ImageNet S (Sketch) Wang et al. (2019) to ensure its performance.
Implementation details. Our experiment is in two parts. In first part we evaluate CMP performance for prompt learning methods. To achieve this, We used pre-trained CLIP (ViT-B16) Radford et al. (2021) model in prompt learning experiments. In second part, we use CLIP (ViT-B16) and (ResNet-50) as backbone while comparing CMP performance with CLIPCalib Murugesan et al. (2025)  [^2] The CLIPCalib prompt learning comparison results are reproduced by using their code published in paper which is available at https://github.com/Bala93/CLIPCalib. in domain generalization. To check the performance of CMP we performed following sets of experiments.
@@
  • CMP integration enhances calibration in domain generalization. To analyze the robustness of CMP, we performed experiment in domain generalization settings. The results are presented in tabular formate along with detail discussion in section 5 and appendix C.
  • CMP improvement against benchmarks. To check the calibration performance of CMP in comparison to benchmarking, we perform detail experiments with two backbone (ResNet-50 and ViT-B16). The results are discussed in section 5 and appendix C.
Due to spatial constraint more experimental detail can be found in appendix C.
Baselines. We compare CMP performance with existing prompt learning methods, such as CoOp Zhou et al. (2022a), CoCoOp Zhou et al. (2022b), ProDA Lu et al. (2022), MaPLe Khattak et al. (2023), and KgCoOp Yao et al. (2023).
Benchmarks. We compared CMP confidence normalization performance against standard CLIP in zero-shot setting, with zero-shot calibration method CLIPCalib Murugesan et al. (2025) and , C-tpt Yoon et al. (2024), B-PEFT Pandey et al. (2024).
Evaluation metrics. To measure the confidence calibration of CMP, we used three evaluation metrics Expected Calibration Error (ECE) Guo et al. (2017), Adaptive Calibration Error (ACE) Guo et al. (2017) and Maximum Calibration Error (MCE) Nixon et al. (2019) and accuracy in our work.
Low-shot settings: We evaluate our method CMP in low-shot settings (1-8 shots per class) to assess the generalization and calibration with limited data. The detailed results are given in appendix Table 6 and Table 7. The results show that CMP maintains robust performance compared to baseline prompt learning methods, in low-shot regime.
Table 1: CMP average calibration performance across 11 datasets for downstream tasks in 8-shot setting. the row with + shows our method combined with existing prompt learning method. The $\downarrow$ represent the smaller value is better, while the $\uparrow$ in accuracy (Acc) denotes the larger value is significant. Bold values shows more significant result.
@@
KgCoOp + (CMP) 76.3 4.11 0.97 4.07
Table 2: Comparison of CMP when combined with prompt learning methods in 8-shot setting. Bold values represent significant performance. In the left column Acc $\uparrow$ demonstrates larger value is significant. In the right column ECE $\downarrow$ shows smaller value is significant.
@@
KgCoOp + (CMP) 72.9 74.3 48.8 52.5 65.4 60.3 2.92 4.13 7.08 13.4 3.08 6.92
Table 3: CMP results in comparison with CLIPCalib Murugesan et al. (2025) with ResNet-50 backbone in 8-shot setting. CMP results for ViT-16 backbone implementation are given in the Appendix C. Bold value represents significance improvement in CMP performance. In ACC part ($\uparrow$ ) denotes larger value better accuracy. while in ECE part ($\downarrow$ ) demonstrate smaller value better calibration.
@@
C-tpt 3.45 3.34 3.55 3.38 3.48 3.47 3.41 3.28 3.50 3.55 3.31 3.35
B-PEFT 3.46 3.36 3.57 3.36 3.58 3.46 3.40 3.32 3.52 3.52 3.38 3.37
CLIPCMP 3.41 3.30 3.52 3.57 3.29 3.36 3.40 3.19 3.46 3.61 3.17 3.31
Table 4: CMP performance results in comparison with CLIPCalib with (ViT-B16) as backbone for domain genralisation in 8-shot setting.
@@ 5 Discussions
In our work, we evaluated the effectiveness of CMP calibration using ECE, ACE, MCE and generalization of CMP using acc metrics. But due to space constraint we reported ECE results in this section. The remaining metrics results can be found in C.
Calibration performance of CLIP-based adaptive models. Figure 1 shows confidence calibration of CLIP in zero-shot settings for CIFAR1010 classification task. The top row (left) in figure shows contrastive loss result without integration of CMP into loss, it has been observed that the confidence distribution for predicting class 0 is spread across all incorrect classes. This indicates overconfidence of the model as the model logits are less focused. The top row (right) shows results of CMP loss for zero-shot tasks, it is shown in figure that the confidence distribution is concentrated near lower values of confidence. This indicates that CMP reduces the overconfidence for incorrect prediction.
@@
The results in Table 1 indicate that integrating CMP consistently reduces ECE, thereby improving model calibration. For example, the baseline CoOp achieves an ECE of $3.16$, while its penalty-enhanced version (CoOp + CMP) reduces ECE to $2.98$, representing a $5.70\%$ improvement. Similarly, the ProDA baseline shows an ECE of $4.13$, which decreases to $3.94$ with the penalty-based version (ProDA + CMP), reflecting a $4.60\%$ reduction.
Notably, MaPLe demonstrates the highest improvement, with ECE decreasing from $3.19$ to $2.88$, corresponding to a $9.70\%$ reduction, highlighting the strong impact of CMP for this specific prompt learning method. In the case of KgCoOp, integrating CMP (KgCoOp + CMP) reduces ECE from $4.24$ to $4.07$, marking a $4.00\%$ improvement.
@@
CMP integration enhances calibration in domain generalization. To evaluate the calibration performance of CMP integration, we analyze the ECE across domain generalization datasets. The results are reported in Table 2. We discuss the improvement in ECE for each prompt learning method while focusing on source and target domain. It is observed in Table 2, CoOp achieve an average ECE of $8.01$ reflecting significant calibration error. By incorporating our penalty to this method the average ECE is reduced to $7.32$ representing 8.568.56% improvement. The ECE improvement shows the effective mitigation of overconfidence in CoOp method, particularly in domain generalizing setting. The reported ECE for CoCoOp without CMP is $6.83$, while penalized CoCoOp ECE is reduced to $6.15$, yielding 9.89.8% improvement in calibration error under challenging target domain.
@@
ProDA baseline achieves 8.108.10% ECE on average, with penalty integration ECE reduced to $7.32$ which is 9.609.60% improvement. The baseline method MaPLe shows an average of $6.48$ ECE, which is reduced to $5.25$ by integration of our penalty, marking on average 19.019.0% performance improvement. KgCoOp shows an average ECE of $7.38$ as baseline, while penalty based KgCoOp reports $6.92$ corresponding 6.206.20% reduction in ECE.
@@
From the results given in Table 2, CMP consistently improves in ECE across domain generalization datasets with improvement within range (6.206.20% - 19.019.0%). The results confirms the effectiveness of CMP in mitigating overconfidence under challenging scenarios of domain generalization settings.
@@
CMP improvement against benchmarks. CMP calibration performance evaluation results in comparison with benchmarking method CLIPCalib, C-tpt, B-PEFT and CLIP (baseline) for vision datasets and domain generalization datasets are given in Table 3 and Table 4 respectively.
@@
In Table 3, it can be observed that CLIPCMP achieves an average ECE of $3.31$ significantly improving upon the baseline average ECE i.e 4.494.49% and 2.202.20% improvement over benchmark CLIPCalib. The calibration improvement is notable on various datasets like Flowers102 ($3.76$ to $3.29$), EuroSAT ($3.42$ to $3.17$), and Aircraft ($3.44$ to $3.19$), where CMP excels in resolving overconfidence. The results are clearly in favor of our method, indicating the importance of CMP integration into CLIP framework leads to better-calibrated models. The consistent reduction in ECE across datasets demonstrates the robustness of our method.
@@
Table 4 denotes results of CMP calibration results in comparison with CLIP (baseline) and CLIPCalib, C-tpt, B-PEFT benchmark on domain generalization datasets.
@@
The average calibration error reduces from $5.01$ to $4.71$ when penalty is plugged-in into CLIP framework, showing an improvement of 5.945.94% as compared to baseline while 2.792.79% over the baseline in challenging datasets. We have noticed that our method consistently outperform the benchmark across domain generalization datasets. We have also reported the result for CMP by changing the backbone of CLIP which are given in appendix C.
@@
Tuning strength of CMP.The regularization strength $\lambda$ is chosen via model selection using line search. Models are evaluated against ECE as the metric, as shown in appendix, Table 9. We report ECE results within range ($0.001$ - $0.950$). We have noticed that ECE score is better for a value of lambdalambda ($\lambda$ = $0.01$). All of our experiment results are reported with same value of $\lambda$. It is notable that small regularization such as this amount is sufficient to enable model confidences to realign themselves with the annotated classification objective. More detail results can be found in Table 9.
@@ 6 Conclusions
Appendix A Appendix
Appendix B Theoretical Guarantees
Proposition B.1.
@@ Appendix B Theoretical Guarantees
Let $P(x,y)=\frac{\exp(s(x,y))}{\sum_{k}\exp(s(x,y_{k}))}$ is softmax probability for correct pair. By def. of CMP:
$\text{CMP}=\frac{P(x,y)}{\sum_{y^{\prime}\neq y,\,P(x,y^{\prime})>P(x,y)}P(x,y^{\prime})},$
@@
$P(x,y^{\prime})\leq 1\quad\forall j,\quad\text{and}\quad\sum_{j}P(x,y^{\prime})=1,$
@@
$P(x,y^{\prime})=\sum_{P(x,y^{\prime})>P(x,y)}P(x,y^{\prime})\implies\text{CMP}=1$
@@
Proposition B.2.
@@
$\mathcal{L}_{\text{CMP}}=\frac{1}{N}\sum_{i=1}^{N}\frac{P(x_{i},y_{i})}{\sum_{y^{\prime}\neq y_{i},\,P(x_{i},y^{\prime})>P(x_{i},y_{i})}P(x_{i},y^{\prime})+\epsilon}.$
@@
$P(x,y)>P(x,y^{\prime})\quad\forall y^{\prime}\neq y,$
@@
$\mathcal{L}_{\text{CMP}}(x,y^{\prime})\approx 0\quad\text{for correctly classified pairs.}$
@@
$P(x,y^{\prime})>P(x,y^{\prime})\quad\exists(y^{\prime}\neq y^{\prime}),$
@@
$\mathcal{L}_{\text{CMP}}=\frac{1}{N}\sum_{i=1}^{N}\frac{P(x_{i},y_{i})}{\sum_{y^{\prime}\neq y_{i},\,P(x_{i},y^{\prime})>P(x_{i},y_{i})}P(x_{i},y^{\prime})+\epsilon}.$
@@
$\frac{\partial\mathcal{L}_{\text{CMP}}}{\partial s(x_{i},y_{i})}=\frac{\lambda}{N}\cdot\frac{\partial P(x_{i},y_{i})}{\partial s(x_{i},y_{i})}\cdot\frac{\sum_{P(x_{i},y^{\prime})>P(x_{i},y_{i})}P(x_{i},y^{\prime})+\delta-P(x_{i},y_{i})}{\left(\sum_{P(x_{i},y^{\prime})>P(x_{i},y_{i})}P(x_{i},y^{\prime})+\delta\right)^{2}}.$
Since $P(x_{i},y_{i})\leq 1$, the gradient is bounded. The term $\delta>0$ here refers to the infinitesimal step size. Furthermore, the gradient decays smoothly as the gap increases between $P(x_{i},y_{i})$ and $\sum_{P(x_{i},y^{\prime})>P(x_{i},y_{i})}P(x_{i},y^{\prime})$, ensuring numerical stability.
@@
Proposition B.4.
Suppose the CMP model is defined with $\epsilon>0$ so that $0\leq\mathrm{CMP}(s(x,y))\leq 1$ for all scores $s(x,y)$, and assume the parameter domain $\Theta$ is such that the gradient norms of the component losses are uniformly bounded on $\Theta$. That is, there exist finite constants $G_{c},G_{\mathrm{CMP}}<\infty$ with
$\sup_{\theta\in\Theta}\|\nabla_{\theta}\mathcal{L}_{c}(\theta)\|\leq G_{c},\qquad\sup_{\theta\in\Theta}\|\nabla_{\theta}\mathcal{L}_{\mathrm{CMP}}(\theta)\|\leq G_{\mathrm{CMP}}.$
Then the combined loss $\mathcal{L}_{\mathrm{final}}(\theta)=\mathcal{L}_{c}(\theta)+\lambda\mathcal{L}_{\mathrm{CMP}}(\theta)$ satisfies
$\mathcal{L}_{\mathrm{final}}(\theta)\geq 0\quad\text{for all }\theta\in\Theta,\qquad\text{and}\qquad\sup_{\theta\in\Theta}\|\nabla_{\theta}\mathcal{L}_{\mathrm{final}}(\theta)\|\leq G_{c}+\lambda G_{\mathrm{CMP}}<\infty.$
Consequently, under these assumptions the gradient norm of the final loss is uniformly bounded over $\Theta$.
Proof.
First, nonnegativity: $\mathcal{L}_{c}(\theta)\geq 0$ by definition (e.g., cross-entropy/contrastive terms), and since $0\leq\mathrm{CMP}\leq 1$ with $\epsilon>0$ we have $\mathcal{L}_{\mathrm{CMP}}(\theta)=-\log(\mathrm{CMP}(s(x,y))+\epsilon)\geq 0$. Hence $\mathcal{L}_{\mathrm{final}}(\theta)\geq 0$.
For the gradient bound, apply the triangle inequality and the uniform bounds: for any $\theta\in\Theta$,
$\|\nabla_{\theta}\mathcal{L}_{\mathrm{final}}(\theta)\|$ $=\|\nabla_{\theta}\mathcal{L}_{c}(\theta)+\lambda\nabla_{\theta}\mathcal{L}_{\mathrm{CMP}}(\theta)\|$
$\leq\|\nabla_{\theta}\mathcal{L}_{c}(\theta)\|+\lambda\|\nabla_{\theta}\mathcal{L}_{\mathrm{CMP}}(\theta)\|\leq G_{c}+\lambda G_{\mathrm{CMP}}.$
Taking the supremum over $\Theta$ yields the stated bound, finishing the proof. ∎
Remark.
The given proposition provides uniform boundedness of the objective and its gradient on $\Theta$, which helps to prevent gradient explosion in practice. It does not guarantee convergence of optimization iterates.
B.1 Relation to Perplexity